OpenStudy (anonymous):

You would like to have $260,000 in 18 years by making regular deposits at the end of each month in an annuity that pays an annual interest rate of 4.5% compounded monthly. How much of the $260,000 comes from interest? In your calculations, do not round until the final answer. Then, round the monthly payment to the nearest dollar.

OpenStudy (amigatour):

\[A=P(1+\frac{ r }{ n })^{r \times t}\]

OpenStudy (anonymous):

I got that! i have $783.44 but i dont know how to find the interest

OpenStudy (tkhunny):

Yea, no. That's for a single payment only. You have regular, periodic payments. You would like to have $260,000 in 18 years by making regular deposits at the end of each month in an annuity that pays an annual interest rate of 4.5% compounded monthly. How much of the $260,000 comes from interest? In your calculations, do not round until the final answer. Then, round the monthly payment to the nearest dollar. P = The regular Payment i = 0.045 -- Annual interest rate to be compounded monthly j = i/12 = 0.00375 -- Monthly interest rate r = 1+j = 1.00375 -- Monthly Accumulation factor 18*12 = 216 There's all the building blocks you need. Now what?

OpenStudy (anonymous):

I just dont know what the question is asking for?

OpenStudy (tkhunny):

It's first asking for P. n = 1 and P = 1000 n*P = 1000 -- Just one payment at the end of the single period. n = 2 and P = 1000 1000 + 1.00375(1000) = 2003.75 n = 3 and P = 1000 1000 + 1.00375(1000) + 1.00375^2[1000] = 3011.2640625 n = 4 and P = 1000 1000 + 1.00375(1000) + 1.00375^2[1000] + 1.00375^3[1000] = 4022.55630273 You need to keep doing this until you get n all the way up to 216. THEN, you have to get the right payment. $1,000 is not the right answer.